OpenStudy (anonymous):

How do I graph a parametric equation? x = 3 sin3t y = 3 cos3t

OpenStudy (anonymous):

I'm using the calculator by graphmatica :) http://www.graphmatica.com/

OpenStudy (jdoe0001):

heeh

OpenStudy (anonymous):

haha hello

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

sure thing!

OpenStudy (luigi0210):

I'm not sure of the math but you can tell this is gonna be a circle with a radius of 3 centered at the origin

OpenStudy (jdoe0001):

\(\large { \begin{cases} x = 3 sin(3t)\to \cfrac{x}{3}=sin(3t) \\ \quad \\ y = 3 cos(3t)\to \cfrac{y}{3}=cos(3t) \end{cases} \\ \quad \\ recall\implies sin^2(\theta)+cos^2(\theta)=1\qquad thus \\ \quad \\ sin^2(3t)+cos^2(3t)=1\implies \left(\cfrac{x}{3}\right)^2+\left(\cfrac{y}{3}\right)^2=1 \\ \quad \\ \cfrac{x^2}{3^2}+\cfrac{y^2}{3^2}=1\implies x^2+y^2=3^2 }\)

OpenStudy (luigi0210):

If you use \(x=sinx\) and \(y=cosx\), square them and add then you get \(x^2+y^2=sin^2 x+cos^2 x\) Which is \(x^2+y^2=1\)

OpenStudy (anonymous):

1=sin^2x+cos^2x